11. The e.m.f. of an element is 6 V. When the external resistance is 1.1 Ohm the current in the circuit is equal to 3 A. Find the potential drop inside the element and its resistance.


1
Expert's answer
2022-02-01T09:57:27-0500

"I=\frac{\varepsilon}{R+r},\implies"

"r=\frac{\varepsilon}I-R=0.9~Ohm,"

"U_r=\varepsilon\frac r{R+r}=2.7~V."